# Domain Generalization for In-Orbit 6D Pose Estimation

### v1

@@ 4.3.2 Multi-Task Learning
As stated in Section 3.2, our domain generalization strategy relies on multi-task learning to force our network to learn more generic features through auxiliary tasks. At inference, the auxiliary tasks are not inferred, so that the multi-task learning strategy does not increase the computational complexity of the network at inference. Our method relies on both spacecraft segmentation and faces segmentation as auxiliary tasks. As pointed out in Table 4, adding an auxiliary task decreases the errors by on Lightbox and on Sunlamp, compared to a network trained on the heatmap-based keypoint regression task only. Furthermore, if the network is trained on both segmentation tasks, the errors are decreased by on Lightbox and on Sunlamp.
@@
4.3.3 Domain Randomization
As explained in Section 3.3, our domain generalization strategy relies on domain randomization to enlarge the training set distribution in order to force the network to learn domain-invariant features through aggressive data augmentation techniques. Those data augmentations consist of Gaussian Noise, Brightness & Contrast, Hide&Seek [48], texture and exposure [49] augmentations, resulting in distorted training images. Figure 5 provides some examples of training images augmented through this domain randomization strategy.
Table 5 presents the average errors achieved by our method trained using different data augmentation strategies. With no data augmentation, the network completely overfits on the synthetic train set and provides average errors of 0.51m and on Lightbox and 1.6m and on Sunlamp. The testing images are so different from the images seen during training that the network has not learnt any robust features and does not generalize at all. Table 5 also provides the average errors achieved by the network when it is trained with a single data augmentation randomly applied to 50% of the images. While the usual augmentations, i.e., brightness & contrast, hide&seek and Gaussian noise, reduce the average errors by 55% on Lightbox and 25% on Sunlamp, our texture-based augmentation achieves a reduction of 75% of the average pose score on both test sets. Interestingly, the exposure augmentation brings little gain, i.e., 13% on Lightbox while decreasing by 75% the pose score on Sunlamp. Combining these augmentations through a RandAugment [55] policy significantly reduces the pose score, i.e., by 89% on Lightbox and 93% on Sunlamp.
Finally, Table 5 also shows the average errors achieved when the network is trained using RandAugment [55] with all data augmentation policies but one. The augmentations that contribute the most to the generalization abilities of the network are the texture and exposure ones. Without either of these techniques, the average errors on Sunlamp increase by 65%. This can be explained from the fact that the exposure augmentation produces artifacts that are close to the ones observed in Sunlamp while the texture augmentation is well suited for this domain randomization task because it preserves the semantics of the image while randomizing its style.
@@ 4.3.4 PEM Architecture
Table 6 presents the average errors achieved by our method considering several variations from the proposed PEM architecture on three test sets, i.e., synthetic, Lightbox and Sunlamp. It shows that even if directly predicting the pose from the keypoints through a MLP works, expanding the dimension of the keypoints before processing them through this MLP improves the performances on all sets. In addition, pre-processing the keypoints through an attention-based encoder further improves the accuracy. Finally, adding a positional embedding to the keypoint embeddings provides additional gains. Table 6 also shows that the use of the continuous 6D representation for regressing rotations significantly increases the accuracy of the network.

### v2

@@ 4.3.2 Multi-Task Learning
As stated in Section˜3.2, our domain generalization strategy relies on multi-task learning to force our network to learn more generic features through auxiliary tasks. At inference, the auxiliary tasks are not inferred, so that the multi-task learning strategy does not increase the computational complexity of the network at inference. Our method relies on both spacecraft segmentation and faces segmentation as auxiliary tasks. As pointed out in Table˜4, adding an auxiliary task decreases the errors by on Lightbox and on Sunlamp, compared to a network trained on the heatmap-based keypoint regression task only. Furthermore, if the network is trained on both segmentation tasks, the errors are decreased by on Lightbox and on Sunlamp.
4.3.3 Domain Randomization
As explained in Section˜3.3, our domain generalization strategy relies on domain randomization to enlarge the training set distribution in order to force the network to learn domain-invariant features through aggressive data augmentation techniques. Those data augmentations consist of Gaussian Noise, Brightness & Contrast, Hide&Seek [57], texture and exposure [58] augmentations, resulting in distorted training images. Figure˜6 provides some examples of training images augmented through this domain randomization strategy.
Table˜5 presents the average errors achieved by our method trained using different data augmentation strategies. With no data augmentation, the network completely overfits on the synthetic train set and provides average errors of 0.51m and on Lightbox and 1.6m and on Sunlamp. The testing images are so different from the images seen during training that the network has not learnt any robust features and does not generalize at all. Table˜5 also provides the average errors achieved by the network when it is trained with a single data augmentation randomly applied to 50% of the images. While the usual augmentations, i.e., brightness & contrast, hide&seek and Gaussian noise, reduce the average errors by 55% on Lightbox and 25% on Sunlamp, our texture-based augmentation achieves a reduction of 75% of the average pose score on both test sets. Interestingly, the exposure augmentation brings little gain, i.e., 13% on Lightbox while it decreases by 75% the pose score on Sunlamp. Combining these augmentations through a RandAugment [64] policy significantly reduces the pose score, i.e., by 89% on Lightbox and 93% on Sunlamp.
Finally, Table˜5 also shows the average errors achieved when the network is trained using RandAugment [64] with all data augmentation policies but one. The augmentations that contribute the most to the generalization abilities of the network are the texture and exposure ones. Without either of these techniques, the average errors on Sunlamp increase by 65%. This can be explained from the fact that the exposure augmentation produces artifacts that are close to the ones observed in Sunlamp while the texture augmentation is well suited for this domain randomization task because it preserves the semantics of the image while randomizing its style.
@@ 4.3.4 PEM Architecture
Table˜6 presents the average errors achieved by our method considering several variations from the proposed PEM architecture on two test sets, i.e., Lightbox and Sunlamp. It shows that even if directly predicting the pose from the keypoints through a MLP works, expanding the dimension of the keypoints before processing them through this MLP improves the performances on all sets. In addition, pre-processing the keypoints through an attention-based encoder further improves the accuracy. Finally, adding a positional embedding to the keypoint embeddings provides additional gains. Table˜6 also shows that the use of the continuous 6D representation for regressing rotations significantly increases the accuracy of the network.